OpenStudy (anonymous):

If f' is continuous on [a,b] show that \[2\int\limits_{a}^{b}f(x)f'(x)dx=[f(b)]^{2}-[f(a)]^{2}\]

OpenStudy (anonymous):

u substitution

myininaya (myininaya):

let u=f(x) du=f'(x) dx

myininaya (myininaya):

if x=b then u=f(b) if x=a then u=f(a)

OpenStudy (anonymous):

\[u=f,du=f',\int ff'=\frac{f^2}{2}\]

myininaya (myininaya):

\[2\int\limits_{f(a)}^{f(b)}u du=u^2|_{f(a)}^{f(b)}\]

OpenStudy (anonymous):

ooh i like the line!

OpenStudy (zarkon):

make the line longer \[2\left.\int\limits_{f(a)}^{f(b)}u du=u^2\right|_{f(a)}^{f(b)}\] ;)

myininaya (myininaya):

nice

OpenStudy (zarkon):

looks a little too long in this example

OpenStudy (anonymous):

alright thanks

OpenStudy (anonymous):

i guess if it really says "show" then you say the following: consider the function \[\frac{f^2(x)}{2}\] it has derivative \[f(x)f'(x)\] so by the fundamental theorem of calculus part you know \[\int_a^b f(x)f'(x)dx = f^2(b)-f^2(a)\]

OpenStudy (anonymous):

that is all you have to say, because it says "show" not "compute the integral. finis

OpenStudy (zarkon):

forgot your 2 :)

OpenStudy (anonymous):

no i didn't

OpenStudy (anonymous):

oh yes i did!

myininaya (myininaya):

satellite how could you forget the 2 i'm disappointed in you

OpenStudy (anonymous):

\[2\int_a^b f(x)f'(x)dx = f^2(b)-f^2(a)\] there , happy?

myininaya (myininaya):

i guess

OpenStudy (anonymous):

don't you have to get up early tomorrow?

myininaya (myininaya):

no school doesn't start till Wednesday omg i have to get up 5:00 every morning starting then only old people get up that early

OpenStudy (anonymous):

good thing for those alarm clocks and electric lights! \ http://www.youtube.com/watch?v=KoaxO1F_xao

OpenStudy (zarkon):

5am...ouch

myininaya (myininaya):

i have to drive an hour to school and be there at 7:00 to teach cal 2

OpenStudy (anonymous):

ho ho ho

OpenStudy (anonymous):

i bet your students love calc in the morning. love it i am sure!

myininaya (myininaya):

i already know they are going to love it

myininaya (myininaya):

jk

OpenStudy (anonymous):

you know what subject i hate most? calc 2

myininaya (myininaya):

they are going to hate it i have a feeling

myininaya (myininaya):

cal 2?

OpenStudy (anonymous):

methods of differentiation stultifying boring plus just showing off

OpenStudy (anonymous):

look i can find the anti derivative of this and of that and of the other. so damned what? useless i hate it

myininaya (myininaya):

i like to do all those proofs for those formulas

OpenStudy (anonymous):

only one of any use is parts. the rest is a pile. reduction formulas, trig sub etc etc what a waste of brain space

myininaya (myininaya):

last time at thought cal 2 on the final i just made them prove a bunch of formulas (or derive them)

myininaya (myininaya):

i taught cal 2*

OpenStudy (anonymous):

truth is if you picked a function out of a hat the probability that you can find the closed form of its antiderivative is zero

myininaya (myininaya):

its so much more fun to see how the formula works then to just say hey find the length of this function given this formula

OpenStudy (anonymous):

what is calc2* do you look at the bottom of the page?

myininaya (myininaya):

what?

OpenStudy (anonymous):

the asterisk what does it mean?

myininaya (myininaya):

oh i was making a correction to what i said

myininaya (myininaya):

at thought cal 2 should have been i taught cal 2